Hello. I hope that someone can help me out with this problem,
the question is three brick layers and four labour workers make 910$ a day. Ten brick layers and five labour workers make 1 850$.
A)find an equation for this problem, i cam up with
3x + 4y = 910
10x + 5y = 1850
i dont know if this is right, but if not can you help me find a better equation
b) Is there an advantage by eliminating one or the other variable first? Explain
c) how much does 1 brick layer make a day?
If you can help me out i would greatly appreciate it thank you so much
kim

Answer by xerxes0212(2)

3x + 4y = 910 ==> eqn.#1
10x + 5y = 1850 ==> eqn.#2
Simplify eqn.#2. Divide all terms by 5. This results to
2x + y = 370
Multiply both sides of this equation by -4, thus resulting to -8x - 4y = -1480.
Add this equation to eqn. #1 to eliminate y and solve for x:.
3x + 4y = 910
-8x - 4y = -1480
==================
-5x = -570
Therefore x = 114. Substitute this value of x in eqn.#1 to find y.
Then the value of y = 142.
One brick layer makes $114 a day, while a labour worker receives $142 a day.